# CD-NGP: A Fast Scalable Continual Representation for Dynamic Scenes

### v5

@@ Abstract
Current methods for novel view synthesis (NVS) in dynamic scenes encounter significant challenges in managing memory consumption, model complexity, training efficiency, and rendering fidelity. Existing offline techniques, while delivering high-quality results, face challenges from substantial memory demands and limited scalability. Conversely, online methods struggle to balance rapid convergence with model compactness. To address these issues, we propose continual dynamic neural graphics primitives (CD-NGP). Our approach leverages a continual learning framework to reduce memory overhead, and it also integrates features from distinct temporal and spatial hash encodings for high rendering quality. Meanwhile, our method employs parameter reuse to achieve high scalability. Additionally, we introduce a novel dataset featuring multi-view, exceptionally long video sequences with substantial rigid and non-rigid motion, which is seldom possessed by existing datasets. We evaluate the reconstruction quality, speed and scalability of our method on both the established public datasets and our exceptionally long video dataset. Notably, our method achieves an $85\%$ reduction in training memory consumption (less than 14GB) compared to offline techniques and significantly lowers streaming bandwidth requirements (less than 0.4MB/frame) relative to other online alternatives. The experimental results on our long video sequences dataset show the superior scalability and reconstruction quality compared to existing state-of-the-art approaches.
@@
The outline of our continual learning pipeline for dynamic scenes. Our method segments input videos into multiple chunks and encodes the chunks by separate spatial and time hash encodings. Subsequently, we train the auxiliary branches incrementally on each chunk and leverage the re-use of parameters in the base branch to ensure high scalability. The spatial hash features could be encoded according to voxel grids, orthogonally projected plane grids or a combination of both.
@@ Introduction
To address the dynamic reconstruction problem, a number of efforts have been made in the offline setting, requiring access to the entire dataset during the training stage. For example, the methods in [36, 28, 34, 33, 15, 29, 44] propose to use MLP-based representation and achieve high-quality results. There are also methods that speed up the optimization by using the tensor factorization ([40, 17, 7]) and voxel representations [16, 45]. The more recent continual-learning-based methods [27, 48] (also mentioned as online methods) can train models frame by frame, and streamable representations [43] enable the model to be loaded on the fly at inference. However, the offline methods require huge host memory e.g., taking more than $100GB$ when reconstructing multi-view dynamic videos of 300 frames, and the online methods are weak in balancing the convergence speed and model compactness. These limitations in scalability makes them impractical for the wide application on long videos.
@@
In this paper, we propose continual dynamic neural graphics primitives (CD-NGP), a fast and scalable representation for reconstructing dynamic scenes, which enables dynamic reconstruction in long videos. As illustrated in CD-NGP: A Fast Scalable Continual Representation for Dynamic Scenes, CD-NGP derives from three key ideas: dividing the video into segments, encoding spatial and temporal features separately, and reusing features across different segments to ensure high scalability: Firstly, we partition dynamic videos into multiple segments according to time stamps $t$ and perform reconstruction using distinct model branches to facilitate continual learning. Each branch of our model includes MLP regressors, spatial hash encoding for spatial coordinates $(x,y,z)$, and temporal hash encoding for the timestamp $t$. Secondly, the spatial and temporal features are encoded separately to avoid feature confusion in different axes. For each model branch, we sample 3D points $(x,y,z)$ according to frame pixels and cached occupancy grids like [31], and encode the points into spatial features using spatial hash grids. We also encode temporal features using hash grids for the separated $t$ axis. These spatial and temporal features and the viewing direction $\mathbf{d}$ are concatenated and input into MLP regressors for rendering and optimization. Thirdly, we leverage the redundancy in the dynamic scenes and fuse the spatial features of different branches to yield high scalability. We set a larger hash table size ($2^{19}$) in the initial branch (base branch), and set a smaller size ($2^{14}$) for the other branches (auxiliary branches). This asymetric design facilitates reusing the features trained in the base branch, thereby reducing storage consumption and boosting rendering quality.
@@
Figure 1: Comparison between different performance indexes of our method and recent representative baselines. The perceptual consistency is measured by $1-$ LPIPS by AlexNet [57, 24], and the speed is the reciprocal of the training time. The compactness and RAM-Efficiency are mapped by $-log$ function from the model size and memory occupation respectively.
@@
We validate the effectiveness of our method through comprehensive experimental results on both the widely used public datasets and a self-collected dataset. As illustrated in Fig. 1, our method reaches a new balance between rendering quality, model size, and training time while remaining memory-friendly. Compared with prevailing offline baselines, our method consumes $85\%$ less host memory at the training phase, $40\%$ less model size at the inference phase, and delivers comparable results. Compared with online baselines, our method achieves compact model size, high rendering quality, and fast convergence simultaneously. The contributions of this paper are three-fold as follows:
@@
  • We propose a novel method combining separate spatial and temporal hash encodings, achieving high scalability and rendering quality in dynamic scenes.
@@
  • We provide a dataset of multi-view exceptionally long video sequences ($>3$ minutes) with large non-rigid motion, which is seldom possessed by existing datasets.
@@ 1.2 Representations of dynamic scenes
There are mainly 3 variants of NeRF for dynamic scenes. The first class is pure MLP-based representation. D-NeRF [36] leverages a large deformation MLP to model the transition and the trajectories of the objects and extends NeRF to simulated dynamic scenes. DyNeRF [28] uses long latent vectors to encode temporal information along with large MLPs to achieve high-fidelity novel view synthesis in real dynamic scenes. Park et al. [34] propose an interpolation-based MLP representation for dynamic scenes and achieve significant acceleration over DyNeRF. The second class is tensor factorization-based representation. HyperReel [1] proposes a network to predict the sample locations and represent dynamic scenes by key-frame-based TensoRF [9] representation to ensure high fidelity and real-time rendering without any custom CUDA kernels. K-Planes [17] and HexPlane [7] factorize the whole 4D field into the product of 2D planes. Though differently implemented, they both achieve high-fidelity results with intermediate model size (200MB) and training time (100 minutes) on the DyNeRF dataset. The third class is the explicit voxel-based representations. TineuVox [16] leverages the deformation MLP in D-NeRF and implements a 3D voxel grid with multi-scale interpolation, and it achieves high fidelity and fast convergence on the D-NeRF dataset. MixVoxels [45] leverages the standard deviation across the time of all pixels to estimate a variation field to decompose the dynamic scene into static and dynamic voxels. Equipped with different representations for static and dynamic objects, MixVoxels achieves high-fidelity results and super-fast convergence (15 minutes) on DyNeRF dataset. Song et al. [43] train Instant-NGP [31] and TensoRF [9] offline and stream the models along the feature channels for dynamic scenes. Concurrent works are leveraging 4D Gaussian point clouds [55, 49] or 3D Gaussians with deformation networks [54] achieving real-time rendering and high fidelity in dynamic scenes.
@@ 1.3 Continual learning of NeRF
Continual learning (online learning) aims to adapt to new tasks while defying catastrophic forgetting (resisting performance downgrade on the previous tasks) [12]. There are mainly 3 approaches to achieving continual learning: replay, regularization, and parameter isolation. The replay-based methods either store the previous data in a buffer and replay them [39], or use the model to generate the pseudo-ground truth of previous tasks while training on new tasks [41]. The regularization-based methods hold the parameters close to the solution of previous tasks [14]. The parameter-isolation methods use different parameters to cope with different tasks. Recent efforts based on generative replay [6, 35, 11] enable continual learning of NeRF in static scenes and achieve high fidelity. StreamRF [27] applies continual NeRF to dynamic scenes by parameter isolation. Their explicit voxel representation enables fast reconstruction (75 minutes for 300 frames) and could be streamed frame by frame. INV [48] uses a shared color MLP and trains different structure MLPs frame by frame to form a continual NeRF representation. It achieves high reconstruction quality in dynamic scenes.
@@ 2.1 Problem definition and preliminaries
Inspired by the previous NeRF variants [36, 28, 17, 7, 45, 31, 38] for dynamic novel view synthesis and continual learning methods [6, 35, 11, 27, 43, 48] in the static scenes, we focus on the dynamic novel view synthesis problem in a multi-view setting. We solve the problem using continual learning to alleviate the huge memory consumption and bandwidth costs. Unlike the existing replay-based continual learning methods for static scenarios [6, 35], we solve the continual learning problem of dynamic novel view synthesis via parameter isolation. This is because the current offline state-of-the-art methods by tensor factorization [7, 17] or explicit voxels [45] are regressions of the optical attributes in 3D space within a limited time instead of the object trajectory and inherent dynamics. Therefore, the replay method is of limited use in our defined scenario: Since new transient processes are constantly arriving, the model of a limited size is not capable of reconstructing unlimited transient effects. Thus, the replay-based method suffers from catastrophic forgetting. The design of our model follow the parameter isolation approach: Let the time stamps be divided into $N_{chunk}$ chunks, the model consists of $N_{chunk}$ branches to represent different chunks of the dynamic scene accordingly.
@@ 2.2 Representation of short dynamic intervals
To reconstruct the individual chunks of the dynamic scenes, we propose a representation of short dynamic intervals, which serves as the structures of the different branches of our model and handle the dynamic scnes in one chunk efficiently. As shown in [46], naive 4D hash grids for dynamic scenes are short in reconstruction quality. To achieve fast reconstruction and compact model size simultaneously, we represent time and spatial features separately and concatenate them to alleviate feature confusion. The spatial hash features can be encoded along 3D voxels [31], 2D planes [17, 21], or both [38]. The time feature is encoded by short hash tables in a per-frame manner. Following previous state-of-the-art methods [30, 31, 16], we apply a two-fold field function for geometry and color individually. The architecture of the field function is thus defined as:
@@ 2.3 Continual dynamic neural graphics primitives (CD-NGP)
As defined in Section 2.1, we split the model into $N_{chunk}$ individual branches to represent the $N_{chunk}$ chunks of the dynamic scene. Consequently, the spatial features encoded by different model branches could be fused to reduce the number of the parameters and improve scalability. This approach, defined as continual dynamic neural graphics primitives (CD-NGP), is designed to handle the dynamic nature of scenes in the online setting. Equipped with the hash encoders and MLP regressors defined in Section 2.2, the $k$-th branch of the model contains spatial hash encoder $\Psi_{k}(\mathbf{x})$, temporal hash encoder $\Gamma_{k}(t)$, and MLPs $f_{\theta_{1k}},f_{\theta_{2k}}$. Please note the dimensions of the features is independent of the sizes of the hash tables: a larger hash table alleviates hash collision at the cost of more storage consumption, but the features encoded by different sizes of hash tables share the same size, i.e. $\Psi_{k}(x)\in\mathbf{R}^{LF}$ holds for all $k$. To alleviate hash collision and reduce model size simultaneously, we apply an asymmetric hash table design: the hash table in the first branch (base branch) is set as $2^{P_{1}}$, and the hash tables in the subsequent branches (auxiliary branches) are set as $2^{P_{2}}$, where $P_{1}$ is much larger than $P_{2}$. The base branch is designed to reconstruct the scenes with initial movements and static objects, and the auxiliary branches are designed to compensate for residuals. The re-use of the features in the base branch ensures high scalability and alleviates the catastrophic forgetting problem. When ($P_{1},P_{2})=(19,14)$, the hash table in the second branch consumes only $1/32$ of the parameters in the first branch. We substitute the $\Psi(\mathbf{x})$ in Eq. 3 with the fused feature to obtain the volume density $\sigma$ at the $k$-th model branch under the continual learning setting, defined as:
@@
where $\Psi_{0}(\mathbf{x})$ denotes the spatial encoder of the base branch, $\Psi_{k}(\mathbf{x})$ denotes the spatial encoder of the current branch. The view-dependent colors are regressed as in Eq. 4 with the color MLP $f_{\theta_{2k}}$. To cope with large motions and complex transient effects in the long time intervals, we assign $T_{episode}$ chunks as an episode. The MLP regressors are re-trained from initialization in the first branch of each episode to adapt to recent changes. We further illustrate the whole process of our method in Algorithm 1, where $\eta_{init}$ and $\eta_{aux}$ denote the numbers of iterations of the initial branch and the auxiliary branches, respectively. $\Psi_{i},\Gamma_{i},\theta_{1}^{(i)},\theta_{2}^{(i)}$ denote the model parameters in the branches.
@@
The scalability of our method derives from the redundancy of dynamic scenes and the flexibility of hash table sizes. As pointed out by [45], most points ($>90\%$ in DyNeRF dataset by their observation) in the dynamic scenes are static. We thus leverage the shared static points across time and represent the first chunk with a large hash table and the subsequent chunks with small hash tables to ensure scalability. Since the spatial hash tables occupy the most parameters in the model, and most chunks are encoded by the subsequent branches, our asymetric hash table design enables continual learning of dynamic scenes and high scalability in streaming scenarios simultaneously. The total bandwidth consumption of our method is $\mathcal{O}(N_{chunk}\cdot 2^{P_{2}}LF+2^{P_{1}}LF)$, while the previous method [43] requires offline training and consumes $\mathcal{O}(\mathbf{K}\cdot 2^{P_{1}}LF)$ bandwidth, where $\mathbf{K}$ ranging from $0.5$ to $16$ is their bandwidth budget controlling the bitrates in streaming scenarios. If the bandwidth is extremely limited, the hash tables in our model could be arranged as a combination of 3D representations and 2D representations like [38] to reduce the bandwidth cost additionally.

### v6

@@ Abstract
Novel view synthesis (NVS) in dynamic scenes faces persistent challenges in memory consumption, model complexity, training efficiency, and rendering quality. Offline methods offer high fidelity but suffer from high memory usage and limited scalability, while online approaches often trade quality for speed and compactness. We propose Continual Dynamic Neural Graphics Primitives (CD-NGP), a continual learning framework that reduces memory overhead and enhances scalability through parameter reuse. To avoid feature interference in dynamic scenes and improve rendering quality, our method combines spatial and temporal hash encodings, which compactly represent scene structures and motion patterns. We also introduce a new dataset comprising multi-view, long-duration ($>1200$ frames) videos with both rigid and non-rigid motion, which is not found in existing benchmarks. CD-NGP is evaluated on public datasets and our long video dataset, demonstrating superior scalability and reconstruction quality. It significantly reduces training memory usage (to $<$14GB) and requires only 0.4MB/frame in streaming bandwidth on DyNeRF—substantially lower than most online baselines.
@@
Overview of our continual learning pipeline for dynamic scenes. Our method segments multi-view input videos into multiple chunks and subsequently reconstructs them using individual model branches. Each branch contains individual spatial hash encodings for $(x,y,z)$, temporal hash encodings for $t$, and MLP regressors for downstream rendering and optimization. Benefiting from the reuse of the larger hash table in the first branch (base branch) and the fusion of features from the subsequent branches (auxiliary branches), our method achieves high scalability and fast convergence. Leveraging the flexible hash encoding structure, it also demonstrates strong adaptability to different encoding types, including 3D voxel grids, 2D plane grids, or hybrid variants.
@@ Introduction
To address the dynamic reconstruction problem, a number of efforts have been made in the offline setting, requiring access to the entire dataset during the training stage. For example, the methods in [39, 28, 37, 36, 15, 29, 49] propose to use MLP-based representation and achieve high-quality results. Some methods speed up the optimization by using the tensor factorization ([45, 17, 7]) and voxel representations [16, 50]. The more recent continual-learning-based methods [27, 53] (also mentioned as online methods) can train models frame by frame, and streamable representations [48] enable the model to be loaded on the fly at inference.
On the one hand, most offline approaches require loading all video frames of the training set into memory, imposing substantial demands on hardware resources (e.g., approximately 100 GB for the DyNeRF dataset [28] with 20 views $\times$ 300 frames). Lazy loading reduces peak memory usage [61], but still requires prohibitive disk space for full video frame caching. Moreover, the overall resource consumption scales linearly with the number of frames ($O(N_{frames})$), which makes it unsuitable for lengthy video sequences. On the other hand, recent online methods [27, 53] fail to simultaneously achieve fast convergence and compact model size, which constrains their applicability in long-term dynamic scenarios. These limitations motivate us to design a more scalable and memory-efficient representation for modeling dynamic scenes over extended durations.
@@
In this paper, we propose continual dynamic neural graphics primitives (CD-NGP), a fast and scalable representation for reconstructing dynamic scenes, which enables dynamic reconstruction in long videos. As illustrated in CD-NGP: A Fast Scalable Continual Representation for Dynamic Scenes, CD-NGP derives from three key ideas: dividing the video into segments, encoding spatial and temporal features separately, and reusing features across different segments to ensure high scalability: Firstly, we partition dynamic videos into multiple segments according to time stamps $t$ and perform reconstruction using distinct model branches to facilitate continual learning. Each branch of our model includes MLP regressors, spatial hash encoding for spatial coordinates $(x,y,z)$, and temporal hash encoding for the timestamp $t$. Secondly, the spatial and temporal features are encoded separately to avoid feature interference in different axes. For each model branch, we sample 3D points $(x,y,z)$ according to frame pixels and cached occupancy grids like [34], and encode the points into spatial features using spatial hash grids. We also encode temporal features using hash grids for the separated $t$ axis. These spatial and temporal features and the viewing direction $\mathbf{d}$ are concatenated and input into MLP regressors for rendering and optimization. Thirdly, we leverage the redundancy in the dynamic scenes and fuse the spatial features of different branches to yield high scalability. We set a larger hash table size ($2^{19}$) in the initial branch (the base branch), and set a smaller size ($2^{14}$) for the other branches (the auxiliary branches). This asymmetric design facilitates reusing the features trained in the base branch, thereby reducing storage consumption and boosting rendering quality.
@@
Figure 1: Comparison of various performance indices between our method and recent representative baselines. Perceptual consistency is measured by $1-$ LPIPS using AlexNet [63, 24], and convergence speed is defined as the reciprocal of training time. Compactness and RAM efficiency are derived by applying a $-\log$ transformation to model size and memory usage, respectively.
@@
We validate the effectiveness of our method through comprehensive experimental results on both widely used public datasets and a self-collected dataset. As illustrated in Fig. 1, our method reaches a new balance between rendering quality, model size, and training time while remaining memory-friendly. Compared with most prevailing offline baselines, our method requires significantly less host memory during training, maintains a smaller model size at inference, and achieves comparable rendering quality by quantitative metrics. In contrast to online baselines, it simultaneously attains compact model size, high rendering fidelity, and fast convergence. The contributions of this paper are three-fold as follows:
@@
  • We propose a novel method that combines separate spatial and temporal hash encodings to efficiently encode both scene geometry and motion patterns. This design improves scalability and rendering quality in dynamic scenes.
@@
  • We provide a dataset of multi-view exceptionally long video sequences ($>3$ minutes) with large non-rigid motion, which is rarely found in existing datasets.
@@ 1.2 Representations of dynamic scenes
There are mainly 3 variants of NeRF for dynamic scenes. The first class is pure MLP-based representation. D-NeRF [39] leverages a large deformation MLP to model the transition and the trajectories of the objects and extends NeRF to simulated dynamic scenes. DyNeRF [28] uses long latent vectors to encode temporal information along with large MLPs to achieve high-fidelity novel view synthesis in real dynamic scenes. Park et al. [37] propose an interpolation-based MLP representation for dynamic scenes and achieve significant acceleration over DyNeRF. The second class is tensor factorization-based representation. HyperReel [1] proposes a network to predict the sample locations and represent dynamic scenes by key-frame-based TensoRF [9] representation to ensure high fidelity and real-time rendering without any custom CUDA kernels. K-Planes [17] and HexPlane [7] factorize the whole 4D field into the product of 2D planes. Though differently implemented, they both achieve high-fidelity results with intermediate model size (200MB) and training time (100 minutes) on the DyNeRF dataset. The third class is the explicit voxel-based representations. TineuVox [16] leverages the deformation MLP in D-NeRF and implements a 3D voxel grid with multi-scale interpolation, and it achieves high fidelity and fast convergence on the D-NeRF dataset. MixVoxels [50] leverages the standard deviation across the time of all pixels to estimate a variation field to decompose the dynamic scene into static and dynamic voxels. Equipped with different representations for static and dynamic objects, MixVoxels achieves high-fidelity results and super-fast convergence (15 minutes) on the DyNeRF dataset. Song et al. [48] train Instant-NGP [34] and TensoRF [9] offline and stream the models along the feature channels for dynamic scenes. 4K4D [58] uses foreground masks to separate the dynamic objects from the static background, and leverages a point cloud-based method to achieve real-time rendering. Concurrent works are leveraging 4D Gaussian point clouds [61, 54] or 3D Gaussians with deformation networks [60] achieving real-time rendering and high fidelity simultaneously.
@@ 1.3 Continual learning of NeRF
Continual learning (online learning) aims to adapt to new tasks while mitigating catastrophic forgetting (resisting performance downgrade on the previous tasks) [12]. There are mainly 3 approaches to achieving continual learning: replay, regularization, and parameter isolation. The replay-based methods either store the previous data in a buffer and replay them [42], or use the model to generate the pseudo-ground truth of previous tasks while training on new tasks [46]. The regularization-based methods hold the parameters close to the solution of previous tasks [14]. The parameter-isolation methods use disjoint parameters to cope with different tasks.
Depending on the application scenarios, the continual learning methods can be divided into two categories: static and dynamic. On the one hand, several recent works leverage MLP [11] or hash grid representations [6, 38] and replay to enable continual learning of NeRF in static scenes. On the other hand, for the dynamic scenes, the continual learning is mainly based on parameter isolation: StreamRF [27] uses parameter isolation of explicit voxel grids to extend continual NeRF to dynamic scenes. Their representation enables fast reconstruction (75 minutes for 300 frames) and can be streamed frame by frame. INV [53] uses a shared color MLP and trains different structure MLPs frame by frame to form a continual NeRF representation. It delivers high rendering quality in dynamic scenes.
@@ 2 Method
Previous online methods for dynamic scenes [27, 53] do not consider the scalability of the model: their model size scales almost linearly with the number of frames, which is of limited use in long video sequences. Our method combines hash representation and parameter isolation for continual learning and achieves high scalability: it not only supports continually reconstructing the dynamic scene but also scales with a significantly lower model size complexity.
@@ 2.1 Problem definition and preliminaries
Inspired by the previous NeRF variants [39, 28, 17, 7, 50, 34, 41] for dynamic novel view synthesis and continual learning methods [6, 38, 11, 27, 48, 53] in the static scenes, we focus on the dynamic novel view synthesis problem in a multi-view setting. We solve the problem using continual learning to alleviate the huge memory consumption and bandwidth costs. Unlike the existing replay-based continual learning methods for static scenarios [6, 38], we solve the continual learning problem of dynamic novel view synthesis via parameter isolation. This is because the current offline state-of-the-art methods by tensor factorization [7, 17] or explicit voxels [50] are regressions of the optical attributes in 3D space within a limited time instead of the object trajectory and inherent dynamics. Therefore, the replay method is of limited use in our defined scenario: Since new transient processes are constantly arriving, the model of a limited size is not capable of reconstructing unlimited transient effects. Thus, the replay-based method suffers from catastrophic forgetting. The design of our model follows the parameter isolation approach: Let the time stamps be divided into $N_{chunk}$ chunks, and the model consists of $N_{chunk}$ branches to represent different chunks of the dynamic scene accordingly.
@@ 2.2 Representation of short dynamic intervals
To reconstruct the individual chunks of the dynamic scenes, we propose a representation of short dynamic intervals, which serve as the structures of the different branches of our model and handle the dynamic scenes in one chunk efficiently. As shown in [51], naive 4D hash grids for dynamic scenes are short in reconstruction quality. To simultaneously achieve fast reconstruction and compact model size, we use hash encodings to represent temporal and spatial features separately, then we concatenate them to alleviate feature interference. The spatial hash encodings can be implemented as hash tables of 3D voxels [34], orthogonal 2D planes [17, 21], or both [41]. The time feature is encoded by short hash tables in a per-frame manner. Following previous state-of-the-art methods [32, 34, 16], we apply a two-fold field function for geometry and color individually. The architecture of the field function is thus defined as:
@@ 2.3 Continual dynamic neural graphics primitives (CD-NGP)
As defined in Section 2.1, we split the model into $N_{chunk}$ individual branches to represent the $N_{chunk}$ chunks of the dynamic scene. Consequently, the spatial features encoded by different model branches could be fused to reduce the number of parameters and improve scalability. This approach, defined as continual dynamic neural graphics primitives (CD-NGP), is designed to handle the dynamic nature of scenes in the online setting. Equipped with the hash encoders and MLP regressors defined in Section 2.2, the $k$-th branch of the model contains spatial hash encoder $\Psi_{k}(\mathbf{x})$, temporal hash encoder $\Gamma_{k}(t)$, and MLPs $f_{\theta_{1k}},f_{\theta_{2k}}$. Please note the dimensions of the features are independent of the sizes of the hash tables: a larger hash table alleviates hash collision at the cost of more storage consumption, but the features encoded by different sizes of hash tables share the same size, i.e. $\Psi_{k}(x)\in\mathbf{R}^{LF}$ holds for all $k$. To alleviate hash collision and reduce model size simultaneously, we apply an asymmetric hash table design: the hash table in the first branch (base branch) is set as $2^{P_{1}}$, and the hash tables in the subsequent branches (auxiliary branches) are set as $2^{P_{2}}$, where $P_{1}$ is much larger than $P_{2}$. The base branch is designed to reconstruct the scenes with initial movements and static objects, and the auxiliary branches are designed to compensate for residuals. The re-use of the features in the base branch ensures high scalability and alleviates the catastrophic forgetting problem. When ($P_{1},P_{2})=(19,14)$, the hash table in the second branch consumes only $1/32$ of the parameters in the first branch. We substitute the $\Psi(\mathbf{x})$ in Eq. 3 with the fused feature to obtain the volume density $\sigma$ at the $k$-th model branch under the continual learning setting, defined as:
@@
where $\Psi_{0}(\mathbf{x})$ denotes the spatial encoder of the base branch, $\Psi_{k}(\mathbf{x})$ denotes the spatial encoder of the current branch. The view-dependent colors are regressed as in Eq. 4 with the color MLP $f_{\theta_{2k}}$.
We further illustrate the entire process of our method in Algorithm 1, where $\eta_{init}$ and $\eta_{aux}$ denote the numbers of iterations of the first branch and the auxiliary branches, respectively. $\Psi_{i},\Gamma_{i},\theta_{1}^{(i)},\theta_{2}^{(i)}$ denote the model parameters in the branches. The model is trained in a continual learning manner: the first branch is trained from scratch to reconstruct the initial chunk of the dynamic scene, while the MLP decoders of the auxiliary branches are initialized from the previous branch to reconstruct subsequent chunks, facilitating faster convergence. To cope with large motions and complex transient effects in long time intervals, we assign $T_{episode}$ chunks as an episode. The MLP regressors are re-trained from initialization in the starting branch of each episode to adapt to recent changes.
@@
The scalability of our method derives from the redundancy of dynamic scenes and the flexibility of hash table sizes. As pointed out by [50], most points ($>90\%$ in the DyNeRF dataset by their observation) in the dynamic scenes are static. We thus leverage the shared static points across time and represent the first chunk with a large hash table and the subsequent chunks with small hash tables to ensure scalability. Since the spatial hash tables occupy the most parameters in the model, and most chunks are encoded by the subsequent branches, our asymmetric hash table design enables continual learning of dynamic scenes and high scalability in streaming scenarios simultaneously. The total bandwidth consumption of our method is $\mathcal{O}(N_{chunk}\cdot 2^{P_{2}}LF+2^{P_{1}}LF)$, while the previous method [48] requires offline training and consumes $\mathcal{O}(\mathbf{K}\cdot 2^{P_{1}}LF)$ bandwidth, where $\mathbf{K}$ ranging from $0.5$ to $16$ is their bandwidth budget controlling the bitrates in streaming scenarios. If the bandwidth is extremely limited, the hash tables in our model could be arranged as a combination of 3D representations and 2D representations like [41] to reduce the bandwidth cost additionally.